Ethan Manufacturing Inc. produces floor mats for automobiles. The owner, Joseph Ethan, has asked you to assist in estimating maintenance costs. Together, you and Joseph determine that the single best cost driver for maintenance costs is machine hours. These data are from the previous fiscal year for maintenance costs and machine hours:

Month Maintenance Costs Machine Hours
1 $ 2,620 1,710
2 2,780 1,790
3 2,930 1,870
4 3,040 1,890
5 3,120 1,920
6 3,090 1,900
7 3,030 1,880
8 2,870 1,860
9 2,640 1,720
10 2,240 1,120
11 2,250 1,320
12 2,470 1,610

Required:

1. What is the cost equation for maintenance costs using the high-low method?

2. Calculate the mean absolute percentage error (MAPE) for the cost equation you developed in requirement 1.

Complete this question by entering your answers in the tabs below.

  • Required 1
  • Required 2

What is the cost equation for maintenance costs using the high-low method? (Round "slope (unit variable cost)" to 2 decimal places. Include outliers in your calculations.)

Maintenance costs = + ( x )
  • Required 2
  • What is the cost equation for maintenance costs using the high-low method? (Round "slope (unit variable cost)" to 2 decimal places. Include outliers in your calculations.)

    Maintenance costs = + ( x )
Answer #1

Answer (1):

Using high low method:

Machine Manitenence Cost Hours High Value 1,920 $3,120 Low Value 1,120 $2,240 Difference 800 $880

Variable cost per machine hour = 880 / 800 = $1.10 per machine hour

Fixed cost = 3120 - 1920*1.10 = $1,008

As such, cost equation is:

Maintenance cost = $1,008 + $1.10* Machine Hour

Answer (2):

MAPE = 5.28%

Working:

Cost equaltion is:

Maintenance cost = $1,008 + $1.10* Machine Hour

We calculate estimated maintenence cost for months 1 to 12 and calculate mean absolute percentage error (MAPE) as follows:

Month Machine Actual Manitenence Estimate based on Absolute Mean % error Hours Cost (A) cost equation (F) [ABS(A - F) / A) 1

The above excel with 'show formula' is as follows:

A B C E Month Machine Hours Actual Manitenence Estimate based on Cost (A) cost equation (F) 1 2 1 1710 3 2 1790 2620 2780 293